OpenStudy (anonymous):

The legs of a right triangle are 4 units and 7 units. What is the length of the hypotenuse PLEASE HELP!!!!I HAVE AN F IN ALGEBRA PLEASE OH PLEASE HELP LOL

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

to find x, we use the pythagorean theorem a^2 + b^2 = c^2 the legs are 'a' and 'b' c is the hypotenuse 

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 4^2 + 7^2 + x^2 16 + 49 = x^2 65 = x^2 x^2 = 65 x = ???

OpenStudy (anonymous):

@dan815

OpenStudy (shadowlegendx):

Do you get it?

OpenStudy (shadowlegendx):

@Redneckk__Lifee

OpenStudy (shadowlegendx):

Jim explained it very well

OpenStudy (anonymous):

Jim give you the answer already, now it square 2 then just square roof 2 the answer! \[\sqrt{65} =\sqrt{65}\] or you can have... 8.12 :v ( 8.0625... bla bla)